# Morita Rigidity for Kleene Algebras

### v2

@@ 1 Introduction
The theory of modules has previously been extended to the context of semirings (for example see [5]), though we spend some time developing the theory of tensor products since these do not appear to have been previously worked out for Kleene modules, and the standard definition in [5] is inadequate for us because it makes all tensor products of modules over additively idempotent semirings (such as Kleene algebras) trivial. The alternative tensor product defined in [8] does have the properties we need, but the author had worked out what turned out to be the same definition as Katsov’s before discovering Katsov’s paper and so has included a full development of tensor products for Kleene algebras rather than simply extending Katsov’s definition for modules over semirings to the case of Kleene modules.
@@ 2 Semiring Preliminaries
For $S$ and $T$ semirings, an $(S,T)$-bimodule is a commutative monoid $\langle M,+,0\rangle$ equipped with both a left $S$-module map and a right $T$-module map, such that these two structures are compatible in the sense that for any $s\in S$, $t\in T$, and $m\in M$, $s(mt)=(sm)t$. As in the case of semirings, when we need to disambiguate in which module an operation is happening, we use subscripts, as in $s\cdot_{N}\alpha(m+_{M}n\cdot_{M}t)$ where $M$ is a right semiring module and $N$ a left semiring module over a semiring $S$. As in the literature on modules over rings, we sometimes use subscripts to the left or right of a symbol denoting a module over a semiring to indicate both which semiring acts on the module and from which side. Thus if $S$, $T$ are semirings, ${{}_{S}}M$ denotes a left module over $S$, $N_{T}$ denotes a right module over $T$, and ${{}_{S}}B_{T}$ indicates an $(S,T)$-bimodule. This notation is particularly helpful when working with direct sums and tensor products, as we shall later.
@@
It is a well-known fact from universal algebra (see [3] for a proof) that if $\equiv$ is a congruence on an algebraic structure $\mathfrak{A}$, then the set of equivalence classes of $\equiv$ carry a structure of the same type as $\mathfrak{A}$, with operations computed by taking the equivalence class of the corresponding operation in $\mathfrak{A}$ performed on representatives. This is the quotient of the algebra $\mathfrak{A}$ by the congruence $\equiv$. Thus for example if $\equiv$ is a congruence on a semiring $S$ and we use brackets to denote $\equiv$-equivalence classes, then in the quotient of $S$ by $\equiv$ we have for $s,t\in S$ that $[s]+[t]=[s+t]$ and $[s][t]=[st]$. It is easy to use the stipulation that $\equiv$ is a congruence to prove that these operations are well-defined.
@@
For ((1)) we compute that
$t\cdot_{{M}^{\mathrm{op}}}(t^{\prime}\cdot_{{M}^{\mathrm{op}}}m)$ $=t\cdot_{{M}^{\mathrm{op}}}(m\cdot_{M}t^{\prime})=(m\cdot_{M}t^{\prime})\cdot_{M}t=m\cdot_{M}(t^{\prime}\cdot_{T}t)=m\cdot_{M}(t\cdot_{{T}^{\mathrm{op}}}t^{\prime})=$
$=(t\cdot_{{T}^{\mathrm{op}}}t^{\prime})\cdot_{{M}^{\mathrm{op}}}m.$
The case of ((2)) is symmetric. For ((3) note that
@@ 3.1 Examples
Reversing order when appropriate handles the case of homomorphisms of into right modules Specifically, if $M$ and $N$ are $(S,K)$-bimodules, then $\operatorname{Hom}(M,N)$ becomes a right $K$-module when endowed with pointwise addition, the constant function zero, and the right action of $K$ defined by $(\alpha\cdot_{H}k)(m)=\alpha(k\cdot_{M}m)$ for $k\in K$, $m\in M$, and $\alpha\in\operatorname{Hom}(M,N)$.
@@
the endomorphisms of $M$. By the preceding example $\mathop{\text{End}}(M)$ is a $(K,K)$-bimodule over the semiring reduct of $K$. Now we verify that it in fact carries the structure of a Kleene $(K,K)$-bimodule. For $k\in K$ and $\alpha\in\mathop{\text{End}}(M)$, if $\alpha k\leq\alpha$, then for every $m\in M$, $\alpha(m)k\leq\alpha(m)$, so $\alpha(m)k^{*}\leq\alpha(m)$ for every $m\in M$ and hence $\alpha k^{*}\leq\alpha$. Consequently $\mathop{\text{End}}(K)$ is a right Kleene module over $K$. A similar argument gives that $\mathop{\text{End}}(K)$ is a left Kleene module over $K$, so it remains to verify that the left and right actions of $K$ commute. This is given by the following computation, where $s,k\in K$, $m\in M$, and $\alpha\in\mathop{\text{End}}(K)$.
$(s\alpha)k](m)=(s\alpha)(km)=\alpha(kms)=(\alpha k)(ms)=[s(\alpha k)](m).$
In the special case of $K^{n}$ viewed as a $({K}^{\mathrm{op}},{K}^{\mathrm{op}})$-bimodule with pointwise scalar multiplication, ${\mathop{\text{End}}(M)}^{\mathrm{op}}$ is the $(K,K)$-bimodule of $n\times n$ matrices over $K$. The reason for taking the opposite is that the left and right actions of $K$ on $M$ are reversed in the definition of $\operatorname{Hom}(M,M)=\mathop{\text{End}}(K)$.
@@
Propositional Dynamic Logic. Fix sets $P$ of action atoms and $T$ of test atoms, respectively. [^4] For details concerning propositional dynamic logic see for instance [13]. Let $K$ be the free Kleene algebra generated by $P$, and $B$ the free Boolean algebra generated by $T$. The Lindenbaum-Tarski algebra $A$ of propositional dynamic logic with action atoms $P$ and test atoms $T$ can be viewed as a $(B,B)$-bimodule via the subalgebra bimodule structure, but also has the structure of a left $K$-module via the action $pa=[\langle p\rangle\phi]$ where $p\in K$, $a\in A$, $a=[\phi]$, and brackets denote the equivalence class of a formula in the Lindenbaum-Tarski algebra. This follows immediately from the axioms for propositional dynamic logic.
@@ 3.2 Tensor products of Kleene modules
Fix Kleene algebras $A$, $B$, $C$. A tensor product is a functor $F\mathrel{:}{{}_{A}\text{Mod}_{B}}\times{{}_{B}\text{Mod}_{C}}\rightarrow{{}_{A}\text{Mod}_{C}}$ such that for each $N\in{{}_{B}\text{Mod}_{C}}$, the functors $M.\mapsto F(M,N)$ and $P\mapsto\operatorname{Hom}(N,P)$ form the left and right parts of an adjunction, respectively.
@@ 4 The Morita Skeleton of Semirings and of Kleene Algebras
The Morita skeleton of semirings is a $1$-category with objects isomorphism classes of semirings, [^5] Scott’s trick [17] may be used to represent these equivalence classes as sets. and for $A$ and $B$ semiringss the morphisms between them are all isomorphism classes of semiring $(A,B)$-bimodules (so this category is not locally small). The identity at an algebra $A$ is ${{}_{A}}A_{A}$, and composition of morphisms is given by the operation on isomorphism classes induced by tensor product (which is clearly well-defined because the tensor product is defined via an adjunction). Semirings $A$ and $B$ are called Morita-equivalent precisely when they are isomorphic in the Morita skeleton of semirings.
@@
Let $A$, $B$ be semirings (Kleene algebras) and $h\mathrel{:}A\rightarrow B$ be a homomorphism. The homomorphism module $E_{h}$ is a semiring (Kleene) $(A,B)$-bimodule whose elements and addition are given by the additive reduct of $B$, with the following multiplication actions on $b\in B$: For $a\in A$, $a\cdot_{E_{h}}b=h(a)\cdot_{B}b$. For $b^{\prime}\in B$, $b\cdot_{E_{h}}b^{\prime}=b\cdot_{B}b^{\prime}$.
@@
Next we verify that the tensor product of homomorphism bimodules corrsponds in the Morita skeleton (of semirings or of Kleene algebras) to the homomorphism module of the composition, so that homomorphism modules do indeed give an extension of the category of semirings (Kleene algebras) and homomorphisms to the Morita skeleton.
@@
Now consider $b\otimes c\in E_{g}\otimes E_{f}$. Define $\psi\mathrel{:}E_{f}\otimes E_{g}\rightarrow E_{g\circ f}$ by $\psi(b\otimes c)=b\cdot{E_{g}}c$, and note that this is well-defined because $b\otimes c=1\otimes b\cdot_{E_{g}}c$ for $b\in B$, $c\in C$. It is clearly additive and commutes with the right action of $C$, so we demonstrate that it commutes with the left action of $A$. Let $a\in A$ and note that
@@
Certainly Morita-equivalent Kleene algebras are Morita equivalent as semirings, and so any Kleene algebra $S$ Morita-equivalent to the algebra $K$ is, up to isomorphism, of the form $eMe$ for some finite-dimensional matrix ring $R$ over $K$ and some full idempotent $e\in M$. It is natural to ask whether Kleene algebras whose semiring reducts are Morita equivalent must be Morita equivalent as Kleene algebras. That is to say, if there is a semimodule Morita isomorphism between Kleene algebras $K$ and $S$, must there also be a Kleene algebra Morita isomorphism?
@@
Let $K$ and $S$ be a Kleene algebras. If the semiring reducts of $K$ and $S$ are Morita-equivalent via semiring bimodules, then in fact $K$ and $S$ are Morita-equivalent via Kleene bimodules..
@@ 5 Conclusion and further directions
At the time of this writing Morita equivalence is a new theoretical tool being introduced into the study of Kleene algebras and related structures, but it does have some intriguing potential for applications to the theory of programming language semantics. One observation is that it is possible to freely change scalars from one Kleene-type algebra $K$ to a Morita equivalent one $S$, because if bimodules ${{}_{K}}M_{S}$ and ${{}_{S}}N_{K}$ witness the Morita equivalence of $K$ and $S$, then for any left module ${{}_{K}}P$ we can reversibly pass to the $S$-module ${{}_{S}}P={{}_{S}}N_{K}\otimes{{}_{K}}P$, the reversibility being witnessed by the fact that
@@
It is also of interest to investigate what additional structure can be added to a Kleene algebra (such as order-meets, to form a Kleene lattice, or order-residuals, to form an action algebra; see [12]) while preserving Morita rigidity in the sense that Morita equivalence of semiring reducts implies Morita equivalence of the full algebras (via bimodules which in some appropriate sense preserve the additional structure). This question was suggested to the author by Jose Meseguer at Logic and Applications 2025 in Dubrovník, Croatia The author has investigated this problem but removed results stated in an earlier version due to subtleties concerning homomorphism modules and tensor products pointed out by a diligent, anonymous referee.

### v3

@@ 1 Introduction
The theory of modules has previously been extended to the context of semirings (for example see [5]), though we spend some time developing the theory of tensor products since these do not appear to have been previously worked out for Kleene modules, and the standard definition in [5] is inadequate for us because it makes all tensor products of modules over additively idempotent semirings (such as Kleene algebras) trivial. The alternative tensor product defined in [8] does have the properties we need, though and independently-developed account of tensor products is preserved here.
@@ 2 Semiring Preliminaries
For $S$ and $T$ semirings, an $(S,T)$-bimodule is a commutative monoid $\langle M,+,0\rangle$ equipped with both a left $S$-module map and a right $T$-module map, such that these two structures are compatible in the sense that for any $s\in S$, $t\in T$, and $m\in M$, $s(mt)=(sm)t$. As in the case of semirings, when we need to disambiguate in which module an operation is happening, we use subscripts, as in $s\cdot_{N}\alpha(m+_{M}n\cdot_{M}t)$ where $M$ is a right semiring module and $N$ a left semiring module over a semiring $S$. As in the literature on modules over rings, we sometimes use subscripts to the left or right of a symbol denoting a module over a semiring to indicate both which semiring acts on the module and from which side. Thus if $S$, $T$ are semirings, ${{}_{S}}M$ denotes a left module over $S$, $N_{T}$ denotes a right module over $T$, and ${{}_{S}}B_{T}$ indicates an $(S,T)$-bimodule. This notation is particularly helpful when working with direct sums and tensor products, as we shall later see.
@@
It is a well-known fact from universal algebra (see [3] for a proof) that if $\equiv$ is a congruence on an algebraic structure $\mathfrak{A}$, then the set of equivalence classes of $\equiv$ carries a structure of the same type as $\mathfrak{A}$, with operations computed by taking the equivalence class of the corresponding operation in $\mathfrak{A}$ performed on representatives. This is the quotient of the algebra $\mathfrak{A}$ by the congruence $\equiv$. Thus for example if $\equiv$ is a congruence on a semiring $S$ and we use brackets to denote $\equiv$-equivalence classes, then in the quotient of $S$ by $\equiv$ we have for $s,t\in S$ that $[s]+[t]=[s+t]$ and $[s][t]=[st]$. It is easy to use the stipulation that $\equiv$ is a congruence to prove that these operations are well-defined.
@@
For (1) we compute that
$t\cdot_{{M}^{\mathrm{op}}}(t^{\prime}\cdot_{{M}^{\mathrm{op}}}m)$ $=t\cdot_{{M}^{\mathrm{op}}}(m\cdot_{M}t^{\prime})=(m\cdot_{M}t^{\prime})\cdot_{M}t=m\cdot_{M}(t^{\prime}\cdot_{T}t)=$
$=m\cdot_{M}(t\cdot_{{T}^{\mathrm{op}}}t^{\prime})=(t\cdot_{{T}^{\mathrm{op}}}t^{\prime})\cdot_{{M}^{\mathrm{op}}}m.$
The case of (2) is symmetric. For (3) note that
@@ 3.1 Examples
Reversing order when appropriate handles the case of homomorphisms of into right modules. Specifically, if $M$ and $N$ are $(S,K)$-bimodules, then $\operatorname{Hom}(M,N)$ becomes a right $K$-module when endowed with pointwise addition, the constant function zero, and the right action of $K$ defined by $(\alpha\cdot_{H}k)(m)=\alpha(k\cdot_{M}m)$ for $k\in K$, $m\in M$, and $\alpha\in\operatorname{Hom}(M,N)$. If $M$ and $N$ are both $(K,S)$-bimodules, then $H=\operatorname{Hom}(M,N)$ with the left and right actions defined above in fact has the structure of a Kleene $(S,K)$-bimodule. The only item in the definition of bimodule which needs to be verified is that the left and right actions commute; this is the content of the following computation for $\alpha\in\operatorname{Hom}(M,N)$, $m\in M$, $s\in S$, and $k\in K$:
$[s\cdot_{H}(\alpha\cdot_{H}k)](m)$ $=(\alpha\cdot_{H}k)(m\cdot_{M}s)=\alpha(k\cdot_{M}(m\cdot_{M}s))=$
$=\alpha((k\cdot_{M}m)\cdot_{M}s)=(s\cdot_{H}\alpha)(k\cdot_{M}m)=[(s\cdot_{H}\alpha)\cdot_{H}k](m).$
@@
the endomorphisms of $M$, and note that this has the structure of a $(K,K)$-bimodule. In the special case of $K^{n}$ viewed as a $({K}^{\mathrm{op}},{K}^{\mathrm{op}})$-bimodule with pointwise scalar multiplication, ${\mathop{\text{End}}(M)}^{\mathrm{op}}$ is the $(K,K)$-bimodule of $n\times n$ matrices over $K$. The reason for taking the opposite is that the left and right actions of $K$ on $M$ are reversed in the definition of $\operatorname{Hom}(M,M)=\mathop{\text{End}}(K)$.
@@
Propositional Dynamic Logic. Fix sets $P$ of action atoms and $T$ of test atoms, respectively. [^4] For details concerning propositional dynamic logic see for instance [13]. Let $K$ be the free Kleene algebra generated by $P$, and $B$ the free Boolean algebra generated by $T$. The Lindenbaum–Tarski algebra $A$ of propositional dynamic logic with action atoms $P$ and test atoms $T$ can be viewed as a $(B,B)$-bimodule via the subalgebra bimodule structure, but also has the structure of a left $K$-module via the action $pa=[\langle p\rangle\phi]$ where $p\in K$, $a\in A$, $a=[\phi]$, and brackets denote the equivalence class of a formula in the Lindenbaum–Tarski algebra. This follows immediately from the axioms for propositional dynamic logic.
@@ 3.2 Tensor products of Kleene modules
Fix Kleene algebras $A$, $B$, $C$. A tensor product is a functor $F\mathrel{:}{{}_{A}\text{Mod}_{B}}\times{{}_{B}\text{Mod}_{C}}\rightarrow{{}_{A}\text{Mod}_{C}}$ such that for each $N\in{{}_{B}\text{Mod}_{C}}$, the functors $M\mapsto F(M,N)$ and $P\mapsto\operatorname{Hom}(N,P)$ form the left and right parts of an adjunction, respectively.
@@ 4 The Morita Skeleton of Semirings and of Kleene Algebras
The Morita skeleton of semirings is a $1$-category with objects isomorphism classes of semirings, [^5] Scott’s trick [17] may be used to represent these equivalence classes as sets. and for $A$ and $B$ semirings the morphisms between them are all isomorphism classes of semiring $(A,B)$-bimodules (so this category is not locally small). The identity at an algebra $A$ is ${{}_{A}}A_{A}$, and composition of morphisms is given by the operation on isomorphism classes induced by tensor product (which is clearly well-defined because the tensor product is defined via an adjunction). Semirings $A$ and $B$ are called Morita-equivalent precisely when they are isomorphic in the Morita skeleton of semirings.
@@
Let $A$, $B$ be semirings (Kleene algebras) and $h\mathrel{:}A\rightarrow B$ be a homomorphism. The homomorphism module $E_{h}$ is a semiring (Kleene) $(A,B)$-bimodule whose elements and addition are given by the additive reduct of $B$, with the following multiplication actions on $b\in B$: For $a\in A$, $a\cdot_{E_{h}}b=h(a)\cdot_{B}b$; and for $b^{\prime}\in B$, $b\cdot_{E_{h}}b^{\prime}=b\cdot_{B}b^{\prime}$.
@@
Next we verify that the tensor product of homomorphism bimodules corresponds in the Morita skeleton (of semirings or of Kleene algebras) to the homomorphism module of the composition, so that homomorphism modules do indeed give an extension of the category of semirings (Kleene algebras) and homomorphisms to the Morita skeleton.
@@
Now consider $b\otimes c\in E_{g}\otimes E_{f}$. Define $\psi\mathrel{:}E_{f}\otimes E_{g}\rightarrow E_{g\circ f}$ by $\psi(b\otimes c)=b\cdot_{E_{g}}c$, and note that this is well-defined because $b\otimes c=1\otimes b\cdot_{E_{g}}c$ for $b\in B$, $c\in C$. It is clearly additive and commutes with the right action of $C$, so we demonstrate that it commutes with the left action of $A$. Let $a\in A$ and note that
@@
Certainly Morita-equivalent Kleene algebras are Morita-equivalent as semirings, and so any Kleene algebra $S$ Morita-equivalent to the algebra $K$ is, up to isomorphism, of the form $eMe$ for some finite-dimensional matrix ring $R$ over $K$ and some full idempotent $e\in M$. It is natural to ask whether Kleene algebras whose semiring reducts are Morita-equivalent must be Morita-equivalent as Kleene algebras. That is to say, if there is a semimodule Morita isomorphism between Kleene algebras $K$ and $S$, must there also be a Kleene algebra Morita isomorphism?
@@
Let $K$ and $S$ be a Kleene algebras. If the semiring reducts of $K$ and $S$ are Morita-equivalent via semiring bimodules, then in fact $K$ and $S$ are Morita-equivalent via Kleene bimodules.
@@ 5 Conclusion and further directions
At the time of this writing Morita equivalence is a new theoretical tool being introduced into the study of Kleene algebras and related structures, but it does have some intriguing potential for applications to the theory of programming language semantics. One observation is that it is possible to freely change scalars from one Kleene algebra $K$ to a Morita-equivalent one $S$, because if bimodules ${{}_{K}}M_{S}$ and ${{}_{S}}N_{K}$ witness the Morita equivalence of $K$ and $S$, then for any left module ${{}_{K}}P$ we can reversibly pass to the $S$-module ${{}_{S}}P={{}_{S}}N_{K}\otimes{{}_{K}}P$, the reversibility being witnessed by the fact that
@@
It is also of interest to investigate what additional structure can be added to a Kleene algebra (such as order-meets, to form a Kleene lattice, or order-residuals, to form an action algebra; see [12]) while preserving Morita rigidity in the sense that Morita equivalence of semiring reducts implies Morita equivalence of the full algebras (via bimodules which in some appropriate sense preserve the additional structure). This question was suggested to the author by Jose Meseguer at Logic and Applications 2025 in Dubrovník, Croatia. The author has investigated this problem but removed results stated in an earlier version due to subtleties concerning homomorphism modules and tensor products pointed out by a diligent, anonymous referee.